### v9

# On Achievable Spectral Efficiency Using Adaptive Transmission Over Terrestrial Coherent FSO Links

Abstract
Terrestrial free-space optical (FSO) communication systems, while designed to operate on large unlicensed optical bandwidths, are power-constrained due to strict eye safety regulations. The channel fluctuation inherent in terrestrial FSO links also limits the received optical power. Consequently, the available signal-to-noise ratio (SNR) per Hz could become limited; this holds for future terrestrial systems based on coherent optical communications. An efficient and adaptive transmission mechanism is thus crucial at the optical transmitter. However, a critical assessment of the impact of adaptive transmission in terrestrial FSO systems has received less attention in the literature. This work studies terrestrial coherent FSO communication systems employing adaptive beam transmission while detection receiver operate under shot noise-limited conditions. Specifically, we propose a novel exact closed-form expression for the average spectral efficiency of a coherent FSO system with optimal adaptive transmissions over the gamma-gamma turbulence channel with pointing errors. More importantly, we provide a detailed assessment of the impact of turbulence and pointing error impairments on the coherent FSO system performance, revealing several novel and counterintuitive insights. In particular, the extensive numerical results help elucidate the intricacies of analyzing these terrestrial FSO systems and clarify a few misconceptions alluded to in recent related literature.
I Introduction
Laser-based free-space optical (FSO) communication has gained significant attention as a viable alternative to conventional radio-frequency (RF) and fiber-optic technologies for delivering efficient backhaul connectivity in outdoor wireless networks. This growing interest is driven by several inherent advantages of FSO systems, including the availability of vast unlicensed bandwidth, ease of deployment, inherent physical-layer security, and low-cost transceiver implementation [1]. Consequently, the scope of FSO-based wireless technologies has broadened considerably, finding applications in a wide range of scenarios such as ad hoc wireless networks for tactical and emergency response situations [1], terrestrial last-mile broadband access [2], deep-space communications [3], inter-satellite links [4], and various civil and military operations [5]. In particular, the demand for ultra-high data rates in next-generation wireless networks, including 5G and 6G systems, has spurred renewed interest in coherent FSO technologies, with recent terrestrial demonstrations achieving net data rates exceeding 400 Gbps [6], [7], and 800 Gbps per channel [8].
Detection schemes in most practical FSO communication systems fall into two main categories:
(i) Intensity modulation with direct detection (IM-DD): In this modulation-detection scheme, information is encoded in the intensity of the transmitted optical beam, effectively utilizing only one degree of freedom. At the receiver, the optical signal is detected directly via its impingement on a photodetector. IM-DD is favored in commercial systems due to its simplicity and low-cost implementation but suffers from limited receiver sensitivity, restricting its use to short-range terrestrial links [9], [10].
(ii) Coherent heterodyne detection (HD): Coherent FSO communication systems exploit both the amplitude and phase of the optical carrier to convey information, enabling two degrees of freedom. In HD, a locally generated optical signal—typically offset in frequency from the received signal—is mixed with the incoming beam prior to photodetection. This detection approach significantly improves receiver sensitivity. Additionally, coherent receivers enable high selectivity via narrowband RF filtering, in contrast to the broader optical filtering in IM-DD based FSO communication systems [10].
Despite their numerous advantages, terrestrial FSO communication systems face two primary practical challenges. First, FSO links are highly sensitive to atmospheric effects: rapid channel fluctuations induced by turbulence, and slower attenuation due to scattering and absorption [11]. Second, since FSO communication relies on line-of-sight (LOS) alignment with highly directional laser beams and photodetectors, it is vulnerable to pointing errors. These errors can arise from various physical disturbances such as wind-induced vibrations, transmitter-receiver misalignment due to motion, thermal expansion, or structural sway of buildings. In practice, the severity of such errors depends strongly on the acquisition, tracking, and pointing (ATP) subsystem, which mitigates beam misalignment but cannot fully eliminate residual jitter [11]. Either or both of these impairments can significantly degrade the performance and reliability of FSO systems [12].
Several statistical models for the small-scale intensity fluctu- ations are available in the FSO communication literature with a strong basis in realistic terrestrial propagation conditions. We mention a select few important and relevant models as follows: scintillation in weak turbulence is well modeled using log-normal distribution [13], [14], while the scintillation statistic in the weak-to-strong turbulence regimes is best described by gamma-gamma (GG) distribution [14]. Pointing errors due to jitter are addressed in [15] and extended to nonzero boresight in [16]. More recently, generalized turbulence models such as Málaga ($\mathcal{M}$[17], double-generalized gamma [18], and Fisher–Snedecor $\mathcal{F}$ [19] have been introduced, offering enhanced flexibility across diverse channel conditions.
Transmit optical power in terrestrial FSO communications is limited by stringent eye and skin safety regulations [20]. In contrast, FSO links benefit from virtually unlimited optical bandwidths (tens to hundreds of gigahertz per wavelength) to support ultra-high-speed communications [12], making the available power per hertz a potential limiting factor. While such low-SNR operating points are rarely encountered today, recent demonstrations of extreme-wideband wavelength-division multiplexing (WDM) coherent FSO systems indicate that distributing a fixed transmit power over many carriers may plausibly lead to lower per-channel SNR [7], [21]. In particular, variability in atmospheric attenuation under haze, fog, or pollution can accentuate this effect. Thus, operation in the low-SNR regime may emerge as a relevant consideration for future ultra-wideband WDM terrestrial FSO systems. In addition, irradiance fluctuations (mentioned earlier) present in terrestrial FSO channels degrade the power collected at the receiver [11], but this degradation can be alleviated through power control. For instance, a simple scheme in which the transmitter activates only when the instantaneous channel quality exceeds a predefined threshold can be effective under low-power budgets. Importantly, typical terrestrial FSO channels exhibit extremely high coherence bandwidths (hundreds of gigahertz or more) [22, 23] and relatively long coherence times on the order of milliseconds [24]. These conditions yield a quasi-static, frequency-flat block-fading channel, which is increasingly exploited by recent field trials [6, 7, 8, 21] for adaptive transmission in terrestrial coherent FSO links. This context motivates the subsequent analysis of spectral efficiency in low-SNR regimes, where turbulence-induced fluctuations and limited per-Hz transmit power, especially in ultra-wideband WDM coherent FSO systems, can meaningfully affect achievable throughput.
In contrast to the recent practical advances mentioned above, adaptive terrestrial FSO communication systems have attracted comparatively limited rigorous theoretical analysis in the literature, as summarized next. Prior works on IM-DD and coherent HD based terrestrial FSO communication systems have examined achievable spectral efficiency (SE) under log-normal and gamma-gamma (GG) turbulence models (e.g., [25, 26, 27, 28, 29, 30], and references therein). More recent studies have extended this analysis to generalized turbulence channels, such as those modeled by the double-generalized gamma and Málaga ($\mathcal{M}$) distributions [31], [32], [33], [34]. Despite these contributions, the critical aspect of beam power adaptation at the optical transmitter has been largely overlooked—except in a few cases, such as [25], [26], and [27]. This omission is likely due to the analytical complexity involved in solving the associated optimization problems. In a closely related work [26],the SE of an HD-based adaptive FSO system over GG turbulence is examined. However, this analysis does not account for the critical influence of pointing errors, which are particularly significant in terrestrial coherent optical communication systems. Furthermore, [32] presents a unified SE expression for IM-DD and HD based FSO communication systems operating under the $\mathcal{M}$ distribution, which subsumes the GG distribution as a special case. However, this analysis assumes a fixed transmit beam power, without adaptation. It is also important to note that several FSO related studies [25], [28], [29], [31], [32] lack clarity in the methodology used to simulate varying turbulence conditions. In terrestrial FSO systems, several channel parameters are interdependent—for example, changes in the FSO link length simultaneously affect turbulence severity level, path loss, and pointing error statistics. A rigorous understanding of these interrelations grounded in physical modeling is crucial for accurately evaluating FSO system performance. Neglecting such dependencies may lead to misleading interpretations and flawed insights.
To address these research gaps in terrestrial FSO research, we will utilize adaptive optical signal transmission on existing terrestrial FSO communication systems with an emphasis on spectral efficiency improvement, which basically relies on exploiting the temporal channel fluctuations due to turbulence and pointing error impairments. Specifically, this work aims to provide in-depth insights into the operation of terrestrial FSO communication systems, and provide a baseline SE performance for future adaptive coherent FSO systems with strong practical constraints. The principal contributions of this paper are as follows:
  • We present a novel and exact closed-form expression for the average spectral efficiency (ASE) of beam power adaptation-based coherent FSO communication systems with HD receivers, accounting for gamma-gamma (GG) turbulence and pointing error impairments. Theorem 1, along with its specialization in Corollary 1 for the case without pointing errors, represents the central analytical contribution of this work.
  • The accuracy of the derived closed-form ASE expressions is firmly established through extensive Monte Carlo simulations, demonstrating consistent agreement over wide and practically relevant ranges of signal-to-noise ratios (SNRs), turbulence strength, and pointing error conditions, as illustrated in Figures 1-2.
  • To provide deeper insight into the influence of turbulence and pointing errors on adaptive FSO system performance, the exact ASE is further examined in the asymptotic regimes of both high and low SNR regimes. Theorem 2 characterizes these limiting behaviors in detail, thereby providing clear insights into the performance gains achievable as well as the penalties incurred under practical terrestrial FSO links.
  • Among the several concrete findings, two particularly noteworthy and counterintuitive results are highlighted: i) at high SNRs, the ASE of the optical channel improves with turbulence when strong pointing error condition is considered (cf. Fig. 2); and ii) at low SNRs, the ASE of the terrestrial FSO channel improves with turbulence (cf. Fig. 5). These phenomena are rigorously characterized and analytically justified.
The remainder of this paper is organized as follows. The optical turbulence and pointing error models adopted for terrestrial FSO channel are presented in Subsection II-A. The relevant noise source and the coherent FSO system model are described in Subsections II-B and II-C, respectively. Theorems on the spectral efficiency of the turbulent optical channel are derived in Section III and their extensive numerical analysis is presented in Section IV. Finally, Section V summarizes the main results and provides concluding remarks along with directions for future work.
II Channel and System Models
II-A Gamma–Gamma Turbulence with Pointing Errors
The overall intensity fluctuation, or fading, $I$, of a transmitted laser beam due to atmospheric turbulence and pointing error is modeled as
$I\,=\,I_{l}I_{\mathrm{a}}I_{\mathrm{p}}$ (1)
where $I_{l}$ represents the path loss component, determined by the exponential Beer-Lambert law for a given link length and weather conditions [35], $I_{\mathrm{a}}$ accounts for turbulence-induced fluctuations, and $I_{\mathrm{p}}$ corresponds to pointing-error-induced fluctuations. In this work, we assume $I_{l}=1$. The statistical models for these independent fading components are described next.
Gamma-Gamma Turbulence: The statistics of the intensity fluctuation $I_{\mathrm{a}}$, which spans the weak-to-strong turbulence regimes, are well modeled by the gamma-gamma distribution:
$f_{I_{\mathrm{a}}}(I_{\mathrm{a}})=\frac{2(ab)^{(a+b)/2}}{\Gamma(a)\Gamma(b)}I_{\mathrm{a}}^{{(a+b)/2-1}}K_{a-b}\left(2\sqrt{abI_{\mathrm{a}}}\right);\ I_{\mathrm{a}}>0$ (2)
where $a$ and $b$ are shape parameters [14], $\Gamma(\cdot)$ is the Gamma function [36, Eq. (8.310.1)], and $K_{v}(\cdot)$ is the modified Bessel function of the second kind and order $v$ [36, Eq. (8.494.1)]. For plane-wave optical radiation at the receiver, the GG shaping parameters depend on atmospheric conditions as
$a\,$ $=\,\bigl[{\exp\bigl({{0.49\sigma_{\mathrm{\scriptscriptstyle R}}^{2}}}{\bigl({1+1.11\sigma_{\mathrm{\scriptscriptstyle R}}^{2.4}}\bigr)}^{-7/6}\bigr)-1}\bigr]^{-1},$ (3)
$b\,$ $=\,\bigl[{\exp\bigl({0.51\sigma_{\mathrm{\scriptscriptstyle R}}^{2}\bigl(1+0.69\sigma_{\mathrm{\scriptscriptstyle R}}^{2.4}\bigr)^{-5/6}}\bigr)-1}\bigr]^{-1},$ (4)
where $\sigma_{{\mathrm{\scriptscriptstyle R}}}^{2}=1.23\,C_{\mathrm{n}}^{2}k_{\mathrm{w}}^{7/6}{L^{11/6}}$ is the Rytov variance determining the atmospheric turbulence strength, $k_{\mathrm{w}}=2\pi/{\lambda_{\mathrm{w}}}$ is the optical wave number, $\lambda_{\mathrm{w}}$} is the wavelength, $L$ is the propagation distance, and $C_{\mathrm{n}}^{2}$ is the index of refraction structure parameter [14].
Pointing Error: Pointing error refers to the misalignment of the laser beam with the optical receiver, caused by i) boresight, the fixed displacement between the beam centroid and the receiver center, and ii) jitter, the random temporal displacement due to factors such as building sway, wind, or vibrations. While careful deployment or fast-tracking transmitters can minimize boresight misalignment, residual jitter remains a concern even with a well-designed ATP subsystem [11]. To account for this effect, the severity of jitter in this work is later categorized into representative regimes reflecting different levels of ATP capability (see Table I in Section IV).
After propagating a distance $L$, the laser beam with waist $w_{\mathrm{\scriptscriptstyle L}}$ is incident on the photodetector with aperture radius $r_{\mathrm{\scriptscriptstyle A}}$ [37]. The received optical beam attenuation due to turbulence-induced spread and pointing error, described by radial displacement $r$ between the beam footprint and the detector center, is given by $I_{\mathrm{p}}(r;L)\approx A_{\mathrm{\scriptscriptstyle 0}}\exp(-2r^{2}/w_{\mathrm{\scriptscriptstyle L_{\mathrm{eq}}}}^{2})$ where $w_{\mathrm{\scriptscriptstyle L_{\mathrm{eq}}}}^{2}=w_{\mathrm{\scriptscriptstyle L}}^{2}\sqrt{\pi}\,\text{erf}(v)/(2v\exp(-v^{2}))$ is the (squared) equivalent beam width with $v=\sqrt{\pi}r_{\mathrm{\scriptscriptstyle A}}/\sqrt{2}w_{\mathrm{\scriptscriptstyle L}}$. The fraction of the collected power at $r=0$ is given by
$A_{\mathrm{\scriptscriptstyle 0}}\,\approx\,\left[\mathrm{erf}(v)\right]^{2},$ (5)
where $\text{erf}(\cdot)$ is the standard error function [36, Eq. (8.250.1)], and $0\leqslant A_{\mathrm{\scriptscriptstyle 0}}\leqslant 1$ [15]. For a coherent beam, the beam radius $w_{\mathrm{\scriptscriptstyle L}}$ is related to the beam waist $w_{\mathrm{\scriptscriptstyle 0}}$ at the optical transmitter’s exit aperture by $w_{\mathrm{\scriptscriptstyle L}}=w_{\mathrm{\scriptscriptstyle 0}}(1+\epsilon(\lambda_{\mathrm{w}}L/\pi w_{\mathrm{\scriptscriptstyle 0}}^{2})^{2})^{1/2}$ where $\epsilon=(1+2w_{\mathrm{\scriptscriptstyle 0}}^{2}/\rho_{\scriptscriptstyle 0}^{2}(L))$ is the global coherence parameter, $\rho_{\scriptscriptstyle 0}(L)=(1.46C_{\mathrm{n}}^{2}k_{\mathrm{w}}^{2}L)^{-3/5}$ is the coherence length in turbulence [38], and the transmit beam waist $w_{\scriptscriptstyle 0}$ is related to the transmitter aperture diameter $D$ as $w_{\mathrm{\scriptscriptstyle 0}}={D}/{\sqrt{2}\pi}$ [39]. Farid and Hranilovic [15] derived the PDF of the irradiance component $I_{\mathrm{p}}$ under the assumptions of i.i.d. zero-mean Gaussian distributions for both horizontal and vertical receiver sways (i.e., zero boresight and identical jitters), expressed as
$\phantom{xxxxxxxxxx}f_{I_{\mathrm{p}}}(I_{\mathrm{p}})\,=\,\frac{\xi^{2}}{A_{\scriptscriptstyle 0}^{\xi^{2}}}I_{\mathrm{p}}^{\xi^{2}-1};\ \ \ 0\leqslant I_{\mathrm{p}}\leqslant A_{\scriptscriptstyle 0}\,$, (6)
where the parameter $\xi\triangleq{w_{\mathrm{\scriptscriptstyle L_{eq}}}}/{2\sigma_{\mathrm{e}}}$ is the ratio of the received equivalent beam waist $w_{\mathrm{\scriptscriptstyle L_{eq}}}$ and the standard deviation of the pointing error displacement (jitter). This model is valid when $w_{\mathrm{\scriptscriptstyle L}}/r_{\mathrm{\scriptscriptstyle A}}\gg 1$ and further that the effects of $\xi$ and $A_{\scriptscriptstyle 0}$ parameters are independent [15].
A recent refinement of the pointing error model is introduced in [40], where multiple alternative parameterizations of $\xi$ and $A_{\scriptscriptstyle 0}$ are proposed to enhance modeling accuracy and computational efficiency (see [40, Fig. 5]). In this work, we hence borrow the modified intensity uniform model from [40], which retains the same mathematical form for the probability distribution as in (6), but with redefined parameters given by
$A_{\mathrm{\scriptscriptstyle 0}}$ $\triangleq 1-\exp(-2r_{\mathrm{\scriptscriptstyle A}}^{2}/w_{\mathrm{\scriptscriptstyle L}}^{2}),$ (7)
$\xi^{2}$ $\triangleq r_{\mathrm{\scriptscriptstyle A}}^{2}/(2\sigma_{\mathrm{e}}^{2}A_{\mathrm{\scriptscriptstyle 0}}),$ (8)
Precisely, this choice demonstrates more than an order-of-magnitude reduction in normalized mean square error (NMSE) compared to the Farid–Hranilovic formulation (see [40, Table 1]), thereby offering a substantially more reliable characterization of pointing jitter effects in terrestrial FSO links.
Composite Fading Distribution: Given the distributions in (2) and (6) of the independent constituents $I_{\mathrm{a}}$ and $I_{\mathrm{p}}$, the PDF of the instantaneous composite irradiance $I=I_{\mathrm{a}}I_{\mathrm{p}}$ is developed to [25, Eq. (8)]
$f_{I}(I)=\frac{ab\xi^{2}}{A_{\mathrm{\scriptscriptstyle 0}}\,\Gamma(a)\Gamma(b)}\,G_{1,3}^{3,0}\left(\frac{abI}{A_{\mathrm{\scriptscriptstyle 0}}}\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}\xi^{2}\\ \xi^{2}-1,\,a-1,\,b-1\\ \end{array}\hskip-5.0pt\right).$ (9)
Remark 1.
For longer terrestrial FSO links, and even for medium-haul links operating under strong turbulence, additional impairments such as boresight misalignment, beam wander, and angle-of-arrival fluctuations become dominant and constitute critical practical challenges. In this work, we focus on turbulence and pointing error effects, an assumption that is well-justified for short-haul links and for medium-haul links under weak turbulence, where higher-order impairments are typically negligible or remain within acceptable limits.
II-B Noise Sources
Two fundamental sources of noise interfere in the process of photodetection in FSO receivers: i) the shot noise due to photodetection itself [41], and ii) the thermal noise originating in the receiver circuitry. The shot noise is due to the information optical radiation and the ambient light incident on the receiver lens, and its power grows linearly with the total incident optical power. Importantly, both ambient shot noise and thermal noise are signal-independent, as they remain statistically uncorrelated with the information-carrying optical signal [42].
The IM-DD photodetector output consists of the desirable photocurrent component mapped one-to-one to the intensity of the incoming (information) optical signal, signal-dependent shot noise, ambient shot noise, and thermal noise currents. The operating mode of an IM-DD receiver is determined by the relative strengths of these noise components.
In a coherent HD receiver, in contrast, an optical local oscillator (OLO) signal is mixed with the received signal prior to photodetection. This introduces an OLO-dependent shot noise current in addition to signal-dependent shot noise, ambient shot noise, and thermal noise currents. By ensuring a large OLO optical power, the OLO shot noise component can be made dominant compared to other noise components; this is in fact considered to be the de facto mode of operation of an (ideal) coherent HD receiver [10][42][43].
Remark on the Statistical Characterization of Shot Noise: When the number of received photons per symbol interval is low-as is typical in inter-satellite and ground-to-satellite optical links-the shot noise is accurately modeled using a Poisson distribution [44]. In contrast, for high photon-count regimes, as encountered in most terrestrial FSO applications, a Gaussian approximation is sufficiently accurate. For a rigorous justification of the Gaussian model for high-intensity shot noise, refer [44][45].
II-C Coherent HD Receiver Model
Let $x_{\mathrm{\scriptscriptstyle HD}}\in\mathbb{C}$ denote the transmitted symbol for the coherent HD system respectively. We remind the reader that the baseband (information) signal in a coherent HD scheme represents a complex optical field, whereas the baseband signal in an IM-DD scheme represents optical intensity. The transmitted beam power is constrained by a fixed average power constraint denoted by $\mathbb{E}\left[|x_{\mathrm{\scriptscriptstyle HD}}|^{2}\right]\leqslant P_{\mathrm{avg}}$. As justified in the introduction, owing to millisecond-scale coherence times and hundreds of gigahertz coherence bandwidths, a frequency-flat block fading model is well-justified for a terrestrial optical fading channel subjected to time-varying irradiance fluctuations: each fading state is assumed to span an identical duration, corresponding to a fixed coherence time, with i.i.d. fading realizations across blocks. The block-wise channel state is considered to be perfectly known at both the transmitter and the receiver at all times. In practice, this state needs to be estimated at the receiver only once at the beginning of each block, and conveyed to the transmitter over an error-free feedback link. Within this framework, we consider a FSO transmitter that performs signal adaptation on a block-by-block basis. Adhering to this understanding, the received signal model for the HD detection-based FSO receiver system is developed next.
The complex baseband symbol in a coherent HD receiver is described by
$y_{\mathrm{\scriptscriptstyle HD}}\,=\,h\,x_{\mathrm{\scriptscriptstyle HD}}\,+\,\,w_{\mathrm{\scriptscriptstyle OLO}}\,+\,w_{\mathrm{\scriptscriptstyle OTH}}$ (10)
where $y_{\mathrm{\scriptscriptstyle HD}}\ \in\ \mathbb{C}$ represents its output, $h\in\mathbb{C}$ the fluctuation in the received optical field, $w_{\mathrm{\scriptscriptstyle OLO}}\sim\mathcal{CN}(0,\sigma^{2}_{\mathrm{\scriptscriptstyle OLO}})$ the OLO shot noise, and $w_{\mathrm{\scriptscriptstyle OTH}}$ the total AWGN from all the sources other than the OLO [13, Ch. 3]. As noted earlier, the OLO shot noise is the only dominant noise of consequence. Hence, conditioned on $h$ or more appropriately on $I:=|h|^{2}$, the received SNR is
$\gamma_{\mathrm{\scriptscriptstyle HD}}\,=\,{I\,|x_{\mathrm{\scriptscriptstyle HD}}|^{2}}/{\sigma_{\mathrm{\scriptscriptstyle OLO}}^{2}}.$ (11)
III Achievable Spectral Efficiency With Adaptive Optical Beam Transmission
In an IM-DD FSO system, we notice that the transmitted signal represents optical intensity, and thus must be non-negative. On the other hand, the $x_{\mathrm{\scriptscriptstyle HD}}$ scheme has no such constraint and can utilize two degrees of freedom. As a result, the classical Shannon limit on spectral efficiency for the AWGN channel applies to the HD scheme, but not to the IM-DD system (see [46]). Conditioned on the received SNR in (11), the achievable SE of the coherent optical channel (in nats/s/Hz) based on HD scheme satisfy
$\mathrm{S}_{\scriptscriptstyle\mathrm{HD}}(\lambda)\,$ $=\,\,\ln\left(1+\dfrac{\lambda P(\lambda)}{N_{\mathrm{hd}}}\right)$ (12)
where $\lambda:=I$, $P(\lambda):=|x_{\mathrm{\scriptscriptstyle HD}}|^{2}$ and $N_{\mathrm{hd}}:=\sigma_{\mathrm{\scriptscriptstyle OLO}}^{2}$. The achievable long-term (over blocks) or average spectral efficiency (ASE) is given by
[…]
where the maximization is performed over all feasible beam adaptations choices of $P$ that satisfy $\mathbb{E}[P(\lambda)]\leqslant P_{\mathrm{avg}}$.
Note that ${P_{\mathrm{avg}}}/{N_{\mathrm{hd}}}$ represents the average transmit power normalized by the average noise power, which will henceforth be termed the transmit SNR and denoted by $\mathrm{SNR}$. The optimal beam power adaptation $P^{\,*}$ maximizing (13) is
$\dfrac{P^{\,*}(\lambda)}{N_{\mathrm{hd}}}\,=\,\left(\dfrac{1}{\mu}-\dfrac{1}{\lambda}\right)^{+}$ (14)
with $\mu$ chosen so that the optical power constraint
$\mathbb{E}\biggl[\left(\dfrac{1}{\mu}-\dfrac{1}{\lambda}\right)^{+}\biggr]\,=\,\mathrm{SNR}$ (15)
is satisfied. Additionally, the parameter $\mu$ serves as a channel cutoff, which simplifies the ASE problem in (13) to
[…]
$\phantom{xxxxx}\underbrace{\int_{0}^{\mu}\!\!\ln(\lambda)f_{\lambda}(\lambda)d\lambda\ }_{=\,I_{2}}\,-\,\underbrace{\int_{\mu}^{\infty}\!\!\ln(\mu)f_{\lambda}(\lambda)d\lambda}_{=\,I_{3}}\,\,\Biggr].$ (17)
In evaluating the integrals $I_{1},\,I_{2}$ and $I_{3}$, we will frequently apply an integral identity involving the Meijer-G function [36, Eq. (9.301)], which is reproduced below (see [47] for additional details).
$\int\!\!x^{j-1}G_{p,q}^{m,n}\left(x\,\middle|\hskip-1.49994pt\begin{array}[]{c}a_{1},\ldots,,a_{p}\\ b_{1},\ldots,,b_{q}\end{array}\hskip-5.0pt\right)dx\,=\phantom{xxxxxxxxxxxxxxxxxxxxxxxxxxxx}$ (20)
$G^{m,\,n+1}_{\!p+1,\,q+1}\hskip-2.5pt\left(x\middle|\hskip-3.50006pt\begin{array}[]{c}1,j\!+a_{1},\ldots,j\!+a_{n},j\!+a_{n+1},\ldots,j\!+a_{p}\\ j\!+b_{1},\ldots,j\!+b_{m},\!0,j\!+b_{m+1},\ldots,j\!+b_{q}\end{array}\hskip-5.0pt\right)\!.$ (23)
III-A Exact ASE With and Without Pointing Errors
Theorem 1.
For the coherent HD optical channel subjected to GG turbulence with pointing errors as described in (9), and with perfect channel state information (CSI) at both ends, the exact ASE in (13) is simplified to
[…]
$\phantom{x}\,\,\,+\,\frac{\xi^{2}}{\Gamma(a)\Gamma(b)}\,G_{3,5}^{3,2}\left(\frac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\,1,\,\xi^{2}+1\\[1.00006pt] \xi^{2},\,a,\,b,\,0,\,0\\ \end{array}\hskip-5.0pt\right)\biggr]$ (26)
where $\psi(\cdot)$ is the Euler’s digamma function, and $\mu$ is determined from $\int_{\mu}^{\infty}(1/\mu-1/\lambda)f_{{\lambda}}(\lambda)d\lambda\,=\,\mathrm{SNR}$.
Proof:
The exact ASE expression (from (III)) is
[…]
Evaluation of Integral $I_{1}$: The integral $I_{1}$ can be viewed as
$I_{1}\,=\,\mathbb{E}\left[\ln\lambda\right].$ (28)
We reformulate (28) by taking expectation on the identity $\frac{d}{dt}\lambda^{t}=\lambda^{t}\ln(\lambda)$ and thereupon evaluating at $t=0$ such that
$\phantom{xxxxxxxxxxxx}\mathbb{E}\left[\lambda^{t}\ln(\lambda)\right]\bigl|_{\,t=0}\,\,=\,\frac{d}{dt}\left(\mathbb{E}\left[\lambda^{t}\right]\right)\biggl|_{\,t=0}$
$\text{and hence}\phantom{xxxxxx}I_{1}\,=\,\frac{d}{dt}\left(\int_{0}^{\infty}\!\!\lambda^{t}f_{\lambda}(\lambda)d\lambda\right)\biggl|_{\,t=0}\,.$ (29)
The $t$-th moment in the RHS of (III-A) is computed as
$\int_{0}^{\infty}\!\!\lambda^{t}f_{\lambda}(\lambda)d\lambda\,\,=$
$\frac{ab\xi^{2}}{A_{\mathrm{\scriptscriptstyle 0}}\Gamma(a)\Gamma(b)}\int_{0}^{\infty}\!\!\lambda^{t}G_{1,3}^{3,0}\left(\frac{ab\lambda}{A_{\mathrm{\scriptscriptstyle 0}}}\biggr\rvert\hskip-2.5pt\begin{array}[]{c}\xi^{2}\\ \xi^{2}\!-1,\!a-1,\!b-1\end{array}\hskip-4.49997pt\right)d\lambda.$ (32)
Applying (20) in (III-A), and making the substitutions:
$\phantom{xx}\lim_{x\to 0}\,G_{2,4}^{3,1}\!\left(x\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}\!+t+1\\ \!\xi^{2}\!+t,\!a\!+t,\!b\!+t,\!0\\ \end{array}\hskip-5.0pt\right)=0,\,\,\,\text{and}$ (35)
$\phantom{xx}\lim_{x\to\infty}\,G_{2,4}^{3,1}\!\left(x\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}\!+t+1\\ \!\xi^{2}\!+t,\!a\!+t,\!b\!+t,\!0\\ \end{array}\hskip-5.0pt\right)=\frac{\Gamma\!\left(a\!+t\right)\!\Gamma\!\left(b\!+t\right)}{\xi^{2}\!+t},$ (38)
we get
$I_{1}\,=\,\frac{\xi^{2}}{\Gamma\!(a)\Gamma\!(b)}\,\frac{d}{dt}\left(\left(\frac{A_{\mathrm{\scriptscriptstyle 0}}}{ab}\right)^{t}\frac{\Gamma\left(a+t\right)\Gamma\left(b+t\right)}{\xi^{2}+t}\right)\Biggl|_{\,t=0}.$ (39)
Upon differentiation w.r.t. $t$ and then evaluating the attained expression for $t=0$, we obtain the final $I_{1}$ expression as
$I_{1}\ =\ \ln\left(\frac{A_{\mathrm{\scriptscriptstyle 0}}}{ab}\right)+\psi(a)+\psi(b)-\left(\frac{1}{\xi^{2}}\right)\,\cdot$ (40)
In the above, we have used $d(\ln\Gamma(x))/dx\,=\,\psi(x)$.
Evaluation of Integral $I_{2}$: The integral part $I_{2}$ is now evaluated as
$I_{2}=\int_{0}^{\mu}\!\!\ln(\lambda)f_{\lambda}(\lambda)d\lambda\phantom{xxxxxxxxxxxxxxxssssssssssssxx}\,$
$=\frac{ab\xi^{2}}{A_{\mathrm{\scriptscriptstyle 0}}\Gamma(a)\Gamma(b)}\int_{0}^{\mu}\ln\left(\lambda\right)G_{1,3}^{3,0}\left(\hskip-1.00006pt\frac{ab\lambda}{A_{\mathrm{\scriptscriptstyle 0}}}\biggr\rvert\hskip-4.49997pt\begin{array}[]{c}\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-5.50003pt\right)\hskip-1.49994ptd\lambda.$ (43)
We apply the integration-by-parts technique treating $\ln(\cdot)$ as the first function and the Meijer-G function as the second to get
$I_{2}\,$ $=\,\biggl[G_{2,4}^{3,1}\left(\dfrac{ab\lambda}{A_{\mathrm{\scriptscriptstyle 0}}}\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0\end{array}\hskip-3.50006pt\right)\times\ln(\lambda)$ (46)
$\,\,\,\,\,\,\,\,\,-G_{3,5}^{3,2}\left(\dfrac{ab\lambda}{A_{\mathrm{\scriptscriptstyle 0}}}\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0,0\end{array}\hskip-3.50006pt\right)\biggr]\,\biggl|_{\,\lambda=0}^{\,\lambda=\mu}\,\,\times\,\,\,\dfrac{\xi^{2}}{\Gamma(a)\Gamma(b)}.$ (49)
With the Meijer-G function limits:
$\phantom{xxxxxx}\lim_{x\to 0}\,\,G_{2,4}^{3,1}\left(x\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0\end{array}\hskip-3.50006pt\right)\,\times\,\ln(x)\,=\,0,\,\,\,\text{and}$ (52)
$\phantom{xxxxxx}\lim_{x\to 0}\,\,G_{3,5}^{3,2}\left(x\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0,0\end{array}\hskip-3.50006pt\right)\,=\,0,$ (55)
the final $I_{2}$ expression is given by
$I_{2}\,$ $=\,\left[G_{2,4}^{3,1}\left(\dfrac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0\end{array}\hskip-3.50006pt\right)\right.\times\,\ln(\mu)$ (58)
$\,\,\,\,\,\,\,\,\,-\left.G_{3,5}^{3,2}\left(\dfrac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0,0\end{array}\hskip-3.50006pt\right)\right]\dfrac{\xi^{2}}{\Gamma(a)\Gamma(b)}\cdot$ (61)
Evaluation of Integral $I_{3}$: The $I_{3}$ component is given as
$I_{3}\,$ $=\,\ln(\mu)\int_{\mu}^{\infty}\!\!f_{\lambda}(\lambda)d\lambda,$ (62)
where the integral is computed as
$\int_{\mu}^{\infty}\!\!f_{\lambda}(\lambda)d\lambda\phantom{xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx}$
= $=\,1-\frac{ab\xi^{2}}{A_{\mathrm{\scriptscriptstyle 0}}\Gamma(a)\Gamma(b)}\int^{\mu}_{0}\!\!G_{1,3}^{3,0}\left(\frac{ab\lambda}{A_{\mathrm{\scriptscriptstyle 0}}}\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-5.0pt\right)d\lambda$ (65)
= $=\,1-\frac{\xi^{2}}{\Gamma(a)\Gamma(b)}\,G_{2,4}^{3,1}\left(\frac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\,\xi^{2}+1\\[1.00006pt] \xi^{2},\,a,\,b,\,0\\ \end{array}\hskip-3.50006pt\right).$ (68)
The last equality is obtained by applying the identity in (20) along with the substitution that
$\lim_{x\to 0}\,\,G_{2,4}^{3,1}\left(x\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\,\xi^{2}+1\\[1.49994pt] \xi^{2},\,a,\,b,\,0\\ \end{array}\hskip-3.50006pt\right)\,=\,0.$
Substituting these $I_{1}$, $I_{2}$ and $I_{3}$ expressions back into (27) and consequent simplification gives the final closed-form expression of the ASE in Theorem 1. ∎
The ASE with dynamic beam power adaptation over the GG turbulence channel (but without pointing error) is a useful baseline for comparison as follows.
Corollary 1.
For the coherent HD optical channel subjected to GG turbulence ‘without pointing error’ as described in (2), and with perfect CSI at both ends, the exact ASE is given by
[…]
$+\frac{1}{\Gamma(a)\Gamma(b)}\,G_{2,4}^{2,2}\left({ab}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,1\\[1.00006pt] a,b,0,0\\ \end{array}\hskip-3.50006pt\right)\biggr]$ (71)
where $\mu$ is solved from $\int_{\mu}^{\infty}(1/\mu-1/\lambda)f_{{\lambda}}(\lambda)d\lambda\,=\,\mathrm{SNR}$.
Proof:
In the limiting case where $A_{\mathrm{\scriptscriptstyle 0}}\to 1$ and $\xi^{2}\to\infty$, the fading distribution converges to the pure $G$ distribution. Substituting these limits into (1) and applying the identity from [36, Eq. (9.31.1)] completes the proof of the theorem. ∎
To the best of the authors’ knowledge, the proposed exact ASE solutions, Theorems 1 and Corollary 1, for the GG turbulence channels with and without pointing error are new.
III-B Immediate Insights
The proposed ASE formula is interesting in obtaining quantitative as well as qualitative insights as follows.
  • At high SNRs, contribution from the Meijer-G based term vanishes while $\mu\to 1/\mathrm{SNR}$. Keeping this in mind, and now comparing the ASE of channels with and without pointing error, the ASE loss due to pointing error can be easily quantified in terms of $A_{\mathrm{\scriptscriptstyle 0}}$ and $\xi^{2}$ parameters.
  • Likewise, the ASE loss due to both turbulence and pointing error at high SNR when compared with no channel impairments (pure AWGN) can also be quantified and analyzed.
  • In a fixed terrestrial FSO link, variation in turbulence can lead to change in the pointing error behavior. In such a scenario, the derived closed-form ASE expression also allows an accurate analysis of the impact of these interlinked channel parameters.
A more in-depth and broader exposition of adaptive transmission benefits and the impact of channel impairments will be presented in Section IV, allowing for a wide range of variations in channel turbulence, pointing error, and SNR.
III-C Low-SNR and High-SNR Asymptotic Expressions
[…]
Lemma 1.
The threshold value $\mu$ associated with the optimal beam power adaptation as described in (14) for the transmitted laser beam satisfies
$\mathrm{(i)}\,\,\,\text{At high SNR:}\,\,\mu\,\approx\,\dfrac{1}{\mathrm{SNR}},$ (72)
$\mathrm{(ii)}\,\,\text{At low SNR:}\,\,\,\mu\,\approx\,\dfrac{A_{\mathrm{\scriptscriptstyle 0}}}{4ab}\ln^{2}\left(\dfrac{1}{\mathrm{SNR}}\right).$ (73)
Proof:
Part (i): Substituting (9) into the average optical power constraint (15), we have
$\dfrac{\mu\mathrm{SNR}}{A}=\int_{\mu}^{\infty}\biggl(1-\dfrac{\mu}{\lambda}\biggr)\,G_{1,3}^{3,0}\left(\frac{ab\lambda}{A_{\mathrm{\scriptscriptstyle 0}}}\biggr\rvert\hskip-3.50006pt\begin{array}[]{c}\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-3.50006pt\right)d\lambda,$
where $A\,:=\,{ab\xi^{2}/(A_{\mathrm{\scriptscriptstyle 0}}\Gamma(a)\Gamma(b)})$. Applying the identity in (20) to the power constraint equation gives
$\dfrac{\mu\mathrm{SNR}}{A}\,=\,\dfrac{1}{A}\,$-\, $\dfrac{A_{\mathrm{\scriptscriptstyle 0}}}{ab}G_{2,4}^{3,1}\left(\frac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\,\xi^{2}+1\\[1.00006pt] \xi^{2},\,a,\,b,\,0\\ \end{array}\hskip-3.50006pt\right)$ (76)
$+\,\,\mu\,$ $G_{2,4}^{3,1}\left(\dfrac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1,0\end{array}\hskip-3.50006pt\right).$ (79)
From the power constraint in (15), it follows that $\mu$ exhibits a strictly monotonic inverse relation with $\mathrm{SNR}$, decreasing as $\mathrm{SNR}$ increases. Notably, at high SNRs, the contributions of the Meijer-G function terms in (76) become negligible due to their vanishing behavior for small arguments. Hence limμ→0    μSNR   =   1. This completes the proof of 1).
[…]
$G_{1,3}^{3,0}\left(z\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}\xi^{2}\\ \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-3.50006pt\right)\,\approx\,\dfrac{\sqrt{\pi}\,z^{(-7+2a+2b)/4}}{e^{2\sqrt{z}}}.$ (82)
Only the dominant term of the series expansion is shown in (82), and substituting it into the LHS of the constraint (15) yields
$\mathrm{SNR}\,$ $\approx\,\mathcal{K}\left(\dfrac{B_{1}(ab\mu/A_{\mathrm{\scriptscriptstyle 0}})}{ab\mu/A_{\mathrm{\scriptscriptstyle 0}}}-B_{2}(ab\mu/A_{\mathrm{\scriptscriptstyle 0}})\right)$ (83)
where $\mathcal{K}:=(\sqrt{\pi}\,ab\,\xi^{2})/(A_{\mathrm{\scriptscriptstyle 0}}\,\Gamma(a)\,\Gamma(b))$ is a positive constant independent of the threshold value, and
$B_{1}(x)\,:=\,\,2^{-n}\,\,\,{\Gamma\bigl(n+1,2\sqrt{x}\bigr)},$ (84)
$B_{2}(x)\,:=\,\,2^{2-n}\,{\Gamma\bigl(n-1,2\sqrt{x}\bigr)},$ (85)
where, in turn, $n:=a+b-\tfrac{5}{2}$, and $\Gamma(\cdot,\cdot)$ denotes the upper incomplete Gamma function, whose large-argument series expansion is given below [36]:
$\Gamma(s,t)\,\approx\,t^{s}e^{-t}\biggl(\,\dfrac{1}{t}\,+\,\dfrac{s-1}{t^{2}}\,+\,o\left(\dfrac{1}{t}\right)^{3}\,\biggr).$ (86)
By retaining the two leading terms in (86), (83) simplifies to
$\mathrm{SNR}\,\approx\,\mathcal{K}\left(\dfrac{ab\mu}{A_{\mathrm{\scriptscriptstyle 0}}}\right)^{\left(\frac{a+b}{2}-\frac{11}{4}\right)}e^{-2\sqrt{\frac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu}}.$ (87)
Taking the logarithm of both sides of (87) and retaining the dominant term on the right-hand side, we obtain
$\ln(\mathrm{SNR})\,\approx\,-2\sqrt{{ab\mu}/{A_{\mathrm{\scriptscriptstyle 0}}}}.$ (88)
Solving (88) for $\mu$, we finally arrive at
$\mu\,\approx\,\dfrac{A_{\mathrm{\scriptscriptstyle 0}}}{4ab}\ln^{2}\left(\frac{1}{\mathrm{SNR}}\right).\vskip-10.00002pt$ (89)
This completes the proof of 2). ∎
Theorem 2.
For the coherent HD optical channel subjected to GG turbulence with pointing errors, the asymptotic ASE at low and high SNR are given by
[…]
where the superscripts $\mathrm{low}$’ and $\mathrm{high}$’ stand for the low-SNR (asymptotically 0) and high-SNR (asymptotically $\infty$) regimes respectively.
Proof:
Part (i): At high SNR, the threshold behaves as $\mu\approx\mathrm{SNR}^{-1}$ (from Lemma 1), while the Meijer’s G function becomes negligible for small input arguments, asymptotically approaching zero as
$\lim_{z\to 0}\,\,\,G_{3,5}^{3,2}\left(z\,\biggr\rvert\,\begin{matrix}1,1,\xi^{2}+1\\ \xi^{2},a,b,0,0\end{matrix}\,\right)\,=\,0.$ (92)
With these substitutions in Theorem 1, (91) is proved.
Part (ii): The low-order series expansion of the Meijer-G function in (1) for large input argument $z$ (asymp. $\infty$) is
$G_{3,5}^{3,2}\left(z\,\biggr\rvert\,\,\begin{matrix}1,1,\xi^{2}+1\\ \!\xi^{2},a,b,0,0\end{matrix}\right)\approx\,$ $\,\dfrac{\sqrt{\pi}\,e^{-2\sqrt{z}}}{z^{(7-2a-2b)/4}}\,+\,\dfrac{\Gamma(a)\,\Gamma(b)}{\xi^{2}}\,\times$ (93)
$\left[\ln(z)-\psi(a)-\psi(b)+\dfrac{1}{\xi^{2}}\right].$ (94)
Substituting the above series expansion in (1), we observe
[…]
Comparing (95) with (87) derived in connection with the lemma 1, we find that
[…]
Finally, substituting $\mu$ from  (89) in (96) completes the proof of the low SNR ASE expansion (90) in Theorem 2. ∎
A few interesting observations from the derived theorems are as follows.
  • By applying $A_{\mathrm{\scriptscriptstyle 0}}\to 1$ and $\xi^{2}\to\infty$, the asymptotic ASEs of the GG turbulence channel ‘without pointing errors’ at low and high SNR are computed as
[…]
  • By comparing (91) and (98), a measure of the ASE degradation due to pointing errors at high SNRs can be found as
    $\biggl[\dfrac{1}{\xi^{2}}\,-\,\ln A_{\mathrm{\scriptscriptstyle 0}}\biggr].$ (99)
  • For the GG turbulence channel, the presence of pointing errors decreases the ASE at low SNRs by $A_{\mathrm{\scriptscriptstyle 0}}$ times.
The derived asymptotic results in (90), (91), (97) and (98) are useful for a comprehensive assessment of the impact of variations of turbulence and pointing error statistics over a wide range.
IV Numerical Results and Discussion
In this section, we present the analytical utility of the derived ASE formulas, supported by extensive numerical results. Previous studies on related topics [25], [28], [29], [31], and [32] have suggested, both directly and indirectly, that a wide range of optical turbulence conditions can be achieved by varying the FSO link length. However, link length variations impact the large-scale path loss factor. Additionally, the laser beam waist expands with propagation distance, impacting the pointing error distribution parameters $\xi$ and $A_{\mathrm{\scriptscriptstyle 0}}$. Notably, these studies did not consider these interdependencies. Understanding the interactions among these channel characteristics is essential, as this work aims to accurately analyze the impact of variations in turbulence and pointing error on the ASE performance of terrestrial FSO systems.
TABLE I: FSO SYSTEM AND CHANNEL SETTINGS
Parameter       Symbol Value
Optical wavelength       $\lambda_{\mathrm{w}}$   1550 nm
Tx. Beam waist       $w_{\mathrm{\scriptscriptstyle 0}}$  0.45 cm
FSO channel length       $L$  300 m
Rx. Aperture radius       $r_{\mathrm{\scriptscriptstyle A}}$  1 cm
Tx./Rx. optics eff.       -  100% (assumed)
Path loss factor       -    Unity (assumed)
 Ch. Setting (a): Gamma-Gamma (GG) fading ‘without’ pointing errorxxxx
Turbulence   Rytov       $G$
strength   variance $\sigma_{{\mathrm{\scriptscriptstyle R}}}^{2}$         distribution parameters
Weak   $0.2$    a=  11.651,b= 10.122a\,=\,\,11.651,\,\,b\,=\,10.122
Moderate   11    a=  4.3939,b= 2.5636a\,=\,\,4.3939,\,\,b\,=\,2.5636
Strong   44    a=  4.3407,b= 1.3088a\,=\,\,4.3407,\,\,b\,=\,1.3088
 Ch. Setting (b): GG fading with ‘mild’ pointing error ($\sigma_{\mathrm{e}}=1$ cm)x
Turbulence    Rytov        Pointing Error
strength    variance $\sigma_{{\mathrm{\scriptscriptstyle R}}}^{2}$       distribution parameters
Weak $0.2$   $\xi\,=\,1.8102,\,A_{\mathrm{\scriptscriptstyle 0}}\,=\,0.1526$
Moderate 11   $\xi\,=\,2.1977,\,A_{\mathrm{\scriptscriptstyle 0}}\,=\,0.1035$
Strong 44   $\xi\,=\,3.5510,\,A_{\mathrm{\scriptscriptstyle 0}}\,=\,0.0397$
 Ch. Setting (c): GG fading with ‘strong’ pointing error ($\sigma_{\mathrm{e}}=4$ cm)x
Turbulence    Rytov        Pointing Error
strength    variance $\sigma_{{\mathrm{\scriptscriptstyle R}}}^{2}$       distribution parameters
Weak $0.2$   $\xi\,=\,0.4525,\,A_{\mathrm{\scriptscriptstyle 0}}\,=\,0.1526$
Moderate 11   $\xi\,=\,0.5494,\,A_{\mathrm{\scriptscriptstyle 0}}\,=\,0.1035$
Strong 44   $\xi\,=\,0.8877,\,A_{\mathrm{\scriptscriptstyle 0}}\,=\,0.0397$
To ensure a fair comparison, the FSO Tx-Rx system parameters (e.g., optical wavelength, Tx. beam waist, Rx. aperture) are kept fixed. In addition, a horizontal terrestrial FSO link of ‘fixed length’ is considered, ensuring that the path loss factor remains unchanged. A horizontal distance of 300 m, representative of short-to-medium haul FSO links, is selected to capture realistic turbulence and pointing error effects while keeping other impairments negligible (see Remark 1). To explicitly account for the acquisition, tracking, and pointing (ATP) subsystem, three representative GG channel settings are considered: (a) No pointing error, corresponding to an ideal or “genie-aided” ATP subsystem that fully eliminates jitter, serving as a turbulence-limited benchmark; (b) Mild pointing error, corresponding to a realistic, well-designed ATP subsystem with small residual jitter; and (c) Strong pointing error, corresponding to a poorly designed or absent ATP subsystem (i.e., no ATP), where jitter dominates. Although such severe jitter levels exceed typical values observed in practical FSO systems employing ATP, they are nevertheless included for theoretical completeness, as they highlight the transition from turbulence-limited to jitter-dominated operation.
These settings cover the practical range of ATP capabilities and provide a unified framework for analyzing turbulence-limited, jitter-limited, and intermediate regimes.
To span a realistic weak-to-strong turbulence range while maintaining constant path loss, we recall that the turbulence strength is governed by the Rytov variance
$\sigma_{{\mathrm{\scriptscriptstyle R}}}^{2}=1.23C_{\mathrm{n}}^{2}k_{\mathrm{w}}^{7/6}{L^{11/6}}$ (100)
which underscores its dependence on the index of refraction structure parameter $C_{\mathrm{n}}^{2}$, as discussed below.
Remark 2.
For near-ground horizontal propagation, the parameter $C_{\mathrm{n}}^{2}$ remains relatively constant and primarily varies with altitude-being stronger near the ground and weaker at higher elevations [11, Sec. 12.2][48, Ch. 2].
A few additional important considerations to be noted before proceeding with the numerical results.
  • Recall from (14) and (15) that $\mathrm{SNR}:=P_{\mathrm{avg}}/N_{\mathrm{hd}}$ is the average transmit SNR in all the derived ASE formulae, and that the horizontal axis in Figures 1-7 is the transmit SNR in log scale, not the received SNR. This enables fair comparison and accurate analysis mentioned earlier. If required, the average received SNR can be computed by scaling the transmit $\mathrm{SNR}$ by the mean channel gain $\mathbb{E}[\lambda]$.
  • All ASE values (bits/sec/Hz is the unit of measurement) are plotted by scaling the derived expressions by ${1}/{(\ln 2)}$.
1) Exact ASE behavior: Based on Theorem 1, the exact ASE for HD-detection based coherent optical channel under GG turbulence, with and without pointing errors, is computed and shown in Figures 1 and 2, respectively. The analytical results closely match Monte Carlo simulation outcomes across a wide SNR range and under various turbulence strengths and pointing jitter conditions.
Figure 1: Average spectral efficiency of the GG turbulence channel without pointing errors under the coherent HD detection scheme. The average received SNR is the same as the transmit SNR since the mean channel gain $\mathbb{E}[\lambda]\,=\,1$.
Figure 2: Average spectral efficiency of the GG turbulence channel with pointing jitter under the coherent HD detection scheme. The received SNR is less than the transmit SNR by the mean channel gain $\mathbb{E}[\lambda]$.
Discussion: Figures 1 and 2 provide the following key observations:
  • As shown in Fig. 1, spectral efficiency deteriorates with increasing turbulence at high SNR. This is expected, as stronger turbulence leads to a heavier concentration of low channel gains in the PDF (cf. Fig. 6).
  • As shown in Fig. 2, a significant drop in spectral efficiency at high SNRs is observed in the presence of pointing errors. This is due to the effective SNR reduction at the receiver by the factor $\mathbb{E}[\lambda]$: milder jitter levels result in a moderate reduction in effective SNR, while severe jitter leads to a significant degradation in system performance.
  • Interestingly, Fig. 2 reveals that, under strong pointing errors, spectral efficiency improves with turbulence in the high-SNR regime: the transition from ASE degradation to improvement initiates at transmit SNR of $30$ dB, and correspondingly to a received SNR of $(30-17.58)\approx 12.42$ dB (since $\mathbb{E}[\lambda]=-17.58$ dB). To explain this counterintuitive behavior, an asymptotic high-SNR analysis is carried out next.
2) ASE behavior at High SNRs: The high-SNR asymptotic ASE behavior in (91), as established in Theorem 2, invites comparison with the ASE of an ideal AWGN channel, which is asymptotically given by $\ln\mathrm{SNR}$. The resulting difference—being strictly negative—quantifies the spectral efficiency loss at high SNRs due to the combined impairments of GG turbulence and pointing jitter, and is given by
$\Delta\overline{\mathrm{S}}_{\mathrm{\scriptscriptstyle HD}}\,\triangleq\,\ln\left(\frac{A_{\mathrm{\scriptscriptstyle 0}}}{ab}\right)\,+\,\psi(a)\,+\,\psi(b)\,-\,\left(\frac{1}{\xi^{2}}\right).$ (101)
Discussion: Fig. 3 depicts the spectral efficiency penalty (101) plotted as a function of atmospheric turbulence strength for the mild and strong pointer error channel settings in Table I. Recall that a fixed-length horizontal FSO link is considered, with a laser beam of waist $w_{\mathrm{\scriptscriptstyle 0}}=4.5$ mm at the transmitter (see Table I). The beam footprint at the receiver is broadened due to beam divergence caused by turbulence over the link length. For the strong pointing error channel setting, the enlarged beam footprint leads to
  • a reduction in collected power at the fixed-aperture receiver (i.e., decreased $A_{\mathrm{\scriptscriptstyle 0}}$); and
  • a possible mitigation of the pointing jitter effect (i.e., increased $\xi$).
Figure 3: The average spectral efficiency loss (in magnitude) at high SNRs as a function of turbulence strength for the GG channel with pointing jitter under the coherent HD detection scheme.
As shown in Fig. 3, for the GG fading with strong pointing error channel setting, the ASE at high SNR increases with turbulence in the practical $0<\sigma_{\mathrm{\scriptscriptstyle R}}^{2}\leq 4$ range. This atypical ASE behavior with turbulence arises because the reduction in the pointing jitter effect more than compensates for the loss in collected power at the receiver. Consequently, the ASE penalty (magnitude) decreases approximately from $10$ to $7.1$ bits/s/Hz, indicating a remarkable improvement of nearly $3.0$ bits/s/Hz in the high SNR regime.
For the GG fading with mild pointing error channel setting, no such trade-off is observed, as increased $\xi$ induces marginal improvement while $A_{\mathrm{\scriptscriptstyle 0}}$ continues to decay significantly, leading overall to a monotonic ASE loss with turbulence. The corresponding high-SNR penalty magnitude, illustrated by the blue curve in Fig. 3, ranges from $3.2$ to $6.1$ bits/s/Hz.
In the GG fading without pointing error, on the other hand, the penalty at high SNRs is described by (see (98))
$\Delta\overline{\mathrm{S}}_{\mathrm{\scriptscriptstyle HD}}\,=\,\psi(a)\,+\,\psi(b)\,-\,\ln\left(ab\right)$ (102)
which corresponds to a relatively small reduction, as can be inferred from Fig. 2 (further qualification in Fig. 4).
Figure 4: Average spectral efficiency of the GG turbulence channel without pointing error under the coherent HD scheme: $A_{\mathrm{\scriptscriptstyle 0}}=1$ for “ideal” system condition whereas $A_{\mathrm{\scriptscriptstyle 0}}$ from Table I are considered as representation for realistic system condition.
Two additional observations can be drawn from the ASE results as follows.
  • A common inconsistency in the literature is the omission of the parameter $A_{\mathrm{\scriptscriptstyle 0}}$ when modeling optical fading without pointing errors. While $A_{\mathrm{\scriptscriptstyle 0}}=1$ is often assumed for simplicity, practical FSO systems may exhibit $A_{\mathrm{\scriptscriptstyle 0}}<1$ due to beam divergence and finite aperture effects. Both cases are realistic in practice, but neglecting $A_{\mathrm{\scriptscriptstyle 0}}<1$, when not physically justified, can lead to an overestimated performance, as illustrated in Fig. 4.
    From Fig. 4, the degradation in ASE due to varying turbulence (weak to strong) under the $A_{\mathrm{\scriptscriptstyle 0}}=1$ condition is modest at only $0.5$ bits/s/Hz. In contrast, when $A_{\mathrm{\scriptscriptstyle 0}}<1$, the ASE degradation due to turbulence increases significantly to 2.6 bits/s/Hz. This contrast highlights the importance of modeling $A_{\mathrm{\scriptscriptstyle 0}}$ accurately. The vertical downward-pointing arrows in Fig. 4 further indicate that the ASE reduction associated with $A_{\mathrm{\scriptscriptstyle 0}}<1$ is more severe under strong turbulence and less pronounced under weak turbulence.
  • For the fixed-length optical channel with a constant FSO system configuration, as considered here, an important observation arises regarding the impact of increasing pointing jitter. Recall that $\sigma_{\mathrm{e}}$ represents the standard deviation of the pointing jitter. At high SNR, the corresponding ASE degradation (in nats/s/Hz) attributable solely to jitter can be obtained directly from (91) and is expressed as
[…]
    where $w_{{\mathrm{\scriptscriptstyle L_{eq}}}}$ denotes the received equivalent beam waist (see just above Eq. (5)). As turbulence intensity increases, $w_{\mathrm{L_{eq}}}$ also increases, and vice versa. Hence, the ASE loss associated with a given increment in pointing jitter exhibits an inverse dependence on turbulence strength: the penalty is reduced under stronger turbulence fluctuations and amplified under weaker fluctuations, as illustrated by the vertical down-arrowed lines in Fig. 2.
Figure 5: Spectral efficiency at low SNRs for the GG turbulence channel without pointing jitter under the HD detection scheme (see Table I for turbulence settings).
3) Exact and asymptotic ASE behavior at Low SNRs: As highlighted in the introduction, understanding the ASE limit of terrestrial FSO links at low SNRs is of practical interest. At low power budgets, power control of the transmitted laser beam becomes critical, requiring efficient exploitation of channel fading.
Discussion: Fig. 5 presents the exact ASE for the GG fading channel without pointing error under the HD detection scheme at low SNRs. Somewhat surprisingly, the low-SNR spectral efficiency improves with increasing turbulence. This agrees with (97), which shows that strengthening GG turbulence from profile $(a_{i},b_{i})$ to $(a_{j},b_{j})$ yields the ASE improvement at sufficiently low SNRs, determined by the ratio of the corresponding pre-factors, which can be expressed as
$\text{Relative pre-factor}\,\approx\,\frac{a_{j}b_{j}}{a_{i}b_{i}}.$ (104)
Figure 6: The PDF of the GG fading gain for the settings described in Table I. Notice the larger probabilities for the higher gains with increasing turbulence strength.
Figure 7: Average spectral efficiency results in the low SNR regime for the GG turbulence channel with mild pointing jitter under the coherent HD detection scheme.
This ASE improvement with turbulence at low SNRs is attributed to the enhanced distribution of higher channel gains with increasing turbulence, as shown in Fig. 6. At low SNR, the transmitter, assuming full CSI, adapts the beam power optimally, allocating more power to stronger channel gains while avoiding power wastage on weaker states. Consequently, power adaptation yields a net ASE gain at low SNRs, as confirmed by the numerical results in Fig. 5. For instance, at a target SNR of 10-10 dB/Hz, the ASE improves from $0.1816$ bits/s/Hz under weak turbulence to $0.2434$ bits/s/Hz under strong turbulence ($34\%$ increase). In a wavelength-division multiplexed (WDM) system with $50$ GHz optical bandwidth per channel, this corresponds to an increase from about $9.08$ Gbps to $12.172$ Gbps per channel for the same transmit power budget. Such per-channel low-SNR conditions are not typical in current deployments but may plausibly arise in future ultra-wideband WDM coherent FSO systems where a fixed eye-safety-limited transmit power must be shared among many carriers, with atmospheric variability (e.g., haze or fog) further accentuating this effect. At a lower target SNR of 15-15 dB/Hz, the ASE increases from about $3.52$ Gbps/channel to $5.42$ Gbps/channel ($53\%$ gain). Moreover, the channel’s ASE per unit SNR-known as the wideband slope [49]-improves as the SNR decreases: under strong turbulence it rises from $2.43$ units at 10-10 dB to $3.43$ units at 15-15 dB ($41\%$ increase), while under weak turbulence it increases from $1.82$ to $2.27$ units ($25\%$ gain). Here, ‘unit’ denotes bits/s/Hz per unit SNR, with SNR expressed in linear ratio.
In contrast, the effect of turbulence on the spectral efficiency of the optical channel with pointing error at low SNRs is more involved. As depicted in Figures 7-8, the spectral efficiency initially shows a slight improvement in the weak-to-moderate turbulence regime, but then begins to degrade as turbulence strength increases from moderate to strong levels.
Figure 8: Average spectral efficiency results in the low SNR regime for the GG turbulence channel with strong pointing jitter under the coherent HD detection scheme.
This unusual behavior can be explained using the asymptotic low-SNR spectral efficiency result in (90) from Theorem 2. The net impact of atmospheric turbulence and pointing error on the optical channel’s ASE at low SNR is captured by the scaling coefficient $A_{\mathrm{\scriptscriptstyle 0}}/4ab$, which we rewrite as
$\textrm{ASE scaling factor}\,\triangleq\,A_{\mathrm{\scriptscriptstyle 0}}\times\dfrac{1}{4ab}.$ (105)
Fig. 9 illustrates the variation of the pre-factor in (105) (red curve) valid for turbulence with pointing error, alongside the factor 1/4ab{1}/{4ab} (blue curve), which applies to pure optical turbulence (see (97)). A comparison of the two curves reveals that the parameter $A_{\mathrm{\scriptscriptstyle 0}}$, representing the fraction of collected power, decreases with turbulence-initially slowly in the weak fluctuation regime and then more rapidly under strong turbulence, while the factor 1/4ab{1}/{4ab} increases. For further insight into how $A_{\mathrm{\scriptscriptstyle 0}}$ varies with turbulence, refer to the $A_{\mathrm{\scriptscriptstyle 0}}$ values in Table I, which provide typical values for weak, moderate, and strong turbulence conditions. In general, except under weak fluctuations, the spectral efficiency of the channel at low SNR with both fading and pointing error tends to degrade with increasing turbulence over a broad range of moderate-to-strong conditions.
Figure 9: Variation of the low-SNR ASE pre-factor for the GG turbulence channel with and without pointing jitters under the coherent HD detection scheme.
Interestingly, the asymptotic result (90) indicates that the influence of pointing jitter strength $\xi^{2}$ on the channel’s ASE vanishes in the low-SNR regime: the transmitter operates only during channel peaks, and it can be verified that the distribution of these peaks is largely insensitive to $\xi^{2}$. The low-SNR ASE characterization for mobile RF channels in [50] reveals two key insights: power adaptation reduces to an asymptotically optimal on–off strategy, and transmitter feedback collapses to a single bit information. These simplifications are expected to carry over to terrestrial coherent optical channels at low SNRs, offering practical benefits for FSO system design.
V Conclusion
In this work, we have analyzed the average spectral efficiency of terrestrial FSO communication systems over gamma-gamma turbulence channels with beam power adaptation incorporated at the optical transmitter side. The analysis also highlighted the detrimental impact on the performance due to pointing errors impairments present in typical FSO links.
The proposed exact solutions capture the impact of fading and pointing error parameters, with particular emphasis on the high and low SNR regimes, and highlight an interplay between $A_{\mathrm{\scriptscriptstyle 0}}$ and $\xi^{2}$ as turbulence and pointing error conditions vary. Under mild or negligible pointing error condition-typical of practical systems with well-designed ATP-the expected degradation of spectral efficiency with increasing turbulence is observed. In contrast, for channels with strong pointing error condition-typical of systems with poorly designed or absent ATP, which may occur in low-cost deployments-a notable improvement in spectral efficiency is observed at high SNRs as turbulence increases from weak conditions, since the reduction in pointing jitter outweighs the loss in received power. Overall, these results provide insights across the full spectrum of ATP capabilities, from practical well-designed systems to theoretical or low-cost configurations with no or poorly performing ATP.
At low SNRs, we have shown that the average spectral efficiency can improve with turbulence due to two main factors: i) the distribution of higher fading gains improves with turbulence, and ii) the transmitter efficiently exploits these higher gains through beam power control. However, in the presence of pointing error, the distribution of higher gains deteriorates significantly with turbulence, resulting in a loss of spectral efficiency at low SNRs. From a system design perspective, operation at low SNRs is particularly relevant for terrestrial coherent FSO links for two reasons: first, the stringent per-hertz optical power limitations plausible in future ultra-wideband, long-haul deployments, and second, the superior energy efficiency of wideband communication systems when operated in the power-limited regime. In this light, the spectral efficiency characterization developed in this work provides timely and practically significant insights into the design of next-generation coherent FSO systems.
As future work, the proposed spectral-efficiency framework can be extended to multi-aperture (MIMO) systems, which we identify as a promising direction for advancing research on next-generation terrestrial coherent FSO systems. Similarly, leveraging the derived ASE limits as a reference for analyzing and optimizing practical coherent FSO implementations represents another valuable avenue for further investigation.